# A is for Absorption: Studying Feature Splitting and Absorption in Sparse Autoencoders

### v5

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, yet our understanding of their internal mechanisms lags behind their capabilities. This gap between performance and interpretability raises concerns about the “black box” nature of these models [30]. The field of mechanistic interpretability aims to address this issue by reverse-engineering the internal algorithms of neural networks and performing causal analysis on them [24].
Recent work theorizes that models represent concepts as linear directions in a high-dimensional space, known as the Linear Representation Hypothesis (LRH) [26, 8]. The model is able to represent far more concepts than it has neurons in its hidden space by allowing these concept directions to overlap slighly, known as superposition [8]. Superposition makes it challenging to directly interpret neurons in an LLM, and requires different techniques to extract interpretable feature directions.
As long as the active features in a given LLM activation are sparse and the underlying features follow the LRH, Sparse Autoencoders (SAEs) should be able to recover the true LLM features despite supersition using sparse dictionary learning [27]. Indeed, SAEs have shown potential in decomposing the dense, polysemantic activations of LLMs into more “interpretable” latent features [5, 2].
However, we show that even if all underlying features are linear and sparsely activating, an SAE will still fail to recover the true underlying features if the features form a hierarchy. Instead, an SAE will learn gerrymandered latents that fail to fire on seemingly arbitrary cases where the latent should fire according to the mainline interpretation of the latent. We refer to this failure mode as feature absorption.
@@
Feature absorption poses an obstacle to the practical application of SAEs since it suggests SAE latents may be inherently unreliable classifiers. This is particularly important for applications where we need confidence that latents are fully tracking behaviors, such as bias or deceptive behavior. Furthermore, techniques which seek to describe circuits in terms of a sparse combination of latents will also be more difficult in the presence of feature absorption [21].
An online explorer for our results can be found at https://absorption.streamlit.app.
@@ 2 Background
A linear probe is a simple linear classifier trained on the hidden activations of a neural network, typically using logistic regression (LR) [1].
@@
A k-sparse probe [12] is a linear probe trained on a sparse subset of $k$ neurons or SAE latents. Training a k-sparse probe first requires selecting the $k$ best neurons or SAE latents that in-aggregate act as a good classifier, and then training a standard linear probe on just those $k$ neurons or latents.
Gurnee et al. [12] proposed several methods of estimating the best $k$ neurons or latents to pick, one of which involves first training a LR probe with a L1 loss term, and selecting the $k$ largest elements by probe weight. When we refer to k-sparse probing in this work, we use this method of selecting $k$ latents.
@@
An SAE consists of an encoder, $W_{enc}$, a decoder, $W_{dec}$, and corresponding biases $b_{enc}$ and $b_{dec}$. The SAE has a nonlinearity, $\sigma$, typically a ReLU (or variant such as JumpReLU [29]). Given input activation, $a$, the SAE computes a hidden representation, $f$, and reconstruction, $\hat{a}$:
@@
In an ablation study we examine the downstream causal effect of an SAE latent by computing how patching its activation to 0 changes a downstream metric (e.g. logit difference). A negative ablation effect means intervening on the SAE latent would lower the metric. We follow the work of Marks et al. [21] and provide the algorithm in Appendix A.4.
@@ 3 Toy models of feature absorption
Our initial setup consists of 4 true features, each randomly initialized into orthogonal directions with a 50 dimensional representation vector and unit norm. Each feature fires with magnitude 1.0. Feature $f_{0}$ fires with probability 0.25, and features $f_{1},f_{2}$, and $f_{3}$ fire with probability 0.05. Each SAE training input is created by sampling from these true features and summing the directions of each firing feature. We train a SAE with 4 latents to match the 4 true features using SAELens [15]. The SAE uses L1 loss with l1 coefficient 3e-5, and learning rate 3e-4. We train on 100M activations.
@@
Figure 2(b) shows the encoder and decoder cosine similarities with the true features in the hierarchical co-occurrence setup. Here, we see a clear example of feature absorption. Latent 0 has learned a perfect representation of feature 0, but the encoder has a hole in its recall. Latent 0 fires if feature 0 is active but not feature 1. This is exactly the sort of gerrymandered feature firing pattern we saw in real SAEs for the starting letter task - the encoder has learned to stop the latent firing on specific cases where it looks like it should be firing. In addition, we see that latent 1, which tracks feature 1, has absorbed the feature 0 direction. This results in latent 1 representing a combination of feature 0 and feature 1. We see that the independently firing features 2 and 3 are untouched - the SAE still learns perfect representations of these features. These results are summarized in Table 3. We explore absorption in more toy settings in Appendix A.3.
@@ 4 Experimental setup
In the above prompt, we extract residual stream activations at the _dog token index. These activations are used both for LR probe training and for applying SAEs. We use a train/test split of 80% / 20%, and evaluate only on the test set of the probes, including when running experiments on SAEs. When applying SAEs, we include the SAE error term [21] to avoid changing model output.
@@
We use the base Gemma-2-2B model for most of our studies, along with the full set of Gemma Scope residual stream SAEs of width 16k and 65k released by Deepmind [19]. We also evaluate absorption on our own SAEs trained on Qwen2 0.5B [32] and Llama 3.2 1B [6].
@@ 5.1 Do SAEs learn latents that track first letter information?
We also tried using k=1 sparse probing [12] to identify SAE latents, and found this gives similar results. Further comparison of using k=1 sparse probing vs encoder cosine similarity to identify latents is explored in Appendix A.7.
@@ 5.2 Why do SAE latents underperform?
Figure 6(a) shows a sample prompt containing a series of tokens that start with “S”, and the activations of top SAE latents by ablation score for these tokens. The main “starts with S” latent, 6510, activates on all these tokens except _short. This SAE also has a token-aligned latent, 1085, which activates on variants of the word “short” (“ short”, “SHORT”, etc…). The Neuronpedia dashboard [20] for latent 1085 is shown in Appendix A.15. For the token _short, the main “starts with S” latent does not activate but the “short” latent activates instead.
@@
We conduct an ablation experiment on the _short token, shown in Figure 7(a), and see that latent 1085 has a dramatically larger ablation effect compared with all other SAE latents. This suggests latent 1085 is causally responsible for the model knowing that _short starts with S.
Is it possible that the probe projection is not the causally important component of latent 1085? We conduct another ablation experiment, except now we remove the probe direction from latent 1085 via projection before ablation. The results of this ablation experiment are shown in Figure 7(b). After removing the probe component from latent 1085, it no longer has a significant ablation effect. Thus we know the probe projection of latent 1085 is responsible for model behavior.
@@ 5.3 Measuring feature splitting and feature absorption
A key phenomenon identified from previous studies of SAEs is feature-splitting [2], where a feature represented in a single latent in a smaller SAE can split into two or more latents in a larger SAE. During our experiments, we found strong evidence of feature-splitting in the Gemma Scope SAEs.
@@
We measure feature splitting using k-sparse probing [12] on SAE activations. If increasing the k-sparse probe from $k$ to $k+1$ causes a significant increase in probe F1 score, then the additional SAE latent provides a meaningful signal, and the combination of these $k+1$ latents is likely a feature split. In the example of the uppercase “L” and lowercase “l” split, a k-sparse probe with $k=2$ trained on both these latents should predict “starts with letter L” much better than either latent on its own. Figure 8(a) shows F1 vs K for letters “L” and “N”. The “L” k-sparse probe shows a significant jump in F1 score moving from k=1 to k=2 corresponding to feature splitting, while the F1 score for the “N” k-sparse probe is relatively constant.
@@
If there are more than 200 false negatives per letter, we randomly pick 200 samples to estimate the number of absorptions. We see absorption rate increases with higher sparsity and higher SAE width. Lower L0 likely pushes the SAE to absorb dense features like spelling information, increasing feature sparsity. Feature absorption rate vs L0 for Gemma Scope SAEs layers 0-17 is shown in Figure 9(b). Absorption rate by letter is shown in Appendix A.14. We also train our own set of standard L1 loss SAEs on the first 8 layers of Qwen2 0.5B [32] and Llama 3.2 1B [6], and TopK SAEs [10] on Llama 3.2 1B. In Figure 9(c) we show that absorption occurs in these SAEs as well.
@@ 6 Related work
Probing methods can extract interpretable information from language models, though this does not guarantee the model uses these representations in its computation, and requires labeled data [7].
Prior work has shown that many human-interpretable concepts in LLM activations are represented as linear directions in activation space, known as the linear representation hypothesis [8, 28]. Li et al. [18] used non-linear probes to recover board representations from a transformer trained on Othello scripts (“OthelloGPT”). Nanda [23] later showed that linear representations were not only recoverable but also editable.
Karvonen et al. [16] developed objective metrics for SAE evaluation using Chess and Othello board states, but does not apply these to SAEs trained on LLM activations. Work by Olah et al. [26], Kissane et al. [17], Templeton et al. [31] noted poor precision/recall of SAE latents compared to known proxies. We extend this by showing how sparsity mediates precision/recall across many Gemma Scope SAEs and offer a possible explanation of low recall due to feature absorption.
Engels et al. [9] investigated SAE errors, finding that not all SAE error is linearly decomposable.
@@
Most existing work on SAE interpretability mainly studies max activating examples [5], which may be misleading. There are more rigorous works which only measure precision [2, 31, 17]. Recent work has briefly explored recall and found it to be worse than expected naively, but this remains poorly understood [26]. We build on this work by evaluating precision / recall on a large number of SAEs, and offer a partial explanation for lower-than-expected recall of SAE latents in the form of feature absorption.
@@
Feature splitting was first described in Bricken et al. [2], which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features. Bussmann et al. [4] find that by training an SAE on the decoder of another SAE, a technique called Meta-SAEs, it is possible to break down a single SAE latent like “Einstein” into subcomponents like “German” and “Physicist” and “starts with E”.
@@ 7 Discussion
The primary goal of future work is to find solutions to feature absorption. We are particularly hopeful that work extending Meta-SAEs [4] may solve or mitigate feature absorption. Another possible solution may be attribution dictionary learning [25]. Finally, structured sparsity techniques such as group lasso [13] or hierarchical sparse coding [14] may also be a promising direction of future work.
@@ Conclusion
We identify a form of feature splitting we call “feature absorption”, where more specific latents “steal credit” from more general ones. Absorption creates an interpretability illusion, where a seemingly interpretable latent has arbitrary false negatives in its mainline interpretation. Lower recall poses problems for using SAEs for high-stakes classification or finding sparse circuits [21], as the number of latents needed to characterize model behavior may be much larger than expected.
@@
This project was produced as part of the LASR Labs research program.
@@ A.1 Glossary of Terms
When inserting a SAE into the computation path of the model, errors in SAE reconstruction will propagate to later parts of the model and can change the model output. We refer to the error as the SAE error term, and corresponds to the difference between the SAE output and the original SAE input activation. Marks et al. [21] introduced the idea of adding this error term back to the SAE output to ensure that the SAE does not change model output.
@@
A platform, https://neuronpedia.org, which hosts feature dashboards for popular SAEs [20].
@@
A phenomenon in SAEs introduced by Bricken et al. [2], where a SAE latent tracks a general feature in a narrow SAE, but splits into multiple more specific SAE latents in a wider SAE. For instance, a latent tracking “starts with L” in a narrow SAE may split into a latent tracking “starts with capital L” and a latent tracking “starts with lowercase L” in a wider SAE.
@@ 1. Preliminaries and Assumptions
Let $\mathcal{D}$ be a dataset. We consider a set of features =$\mathcal{F}=\{f_{1},f_{2},\ldots,f_{d}\}$. Each feature $f_{i}\in\mathbb{R}^{k}$ is a vector with unit norm, $\|f_{i}\|_{2}=1$. Features are mutually orthogonal: $f_{i}\cdot f_{j}=\delta_{ij}$ (Kronecker delta), where $\delta_{ij}=1$ if $i=j$ and 00 if $i\neq j$. An activation $h\in\mathbb{R}^{k}$ in the model’s residual stream is a linear combination of active features: $h=\sum_{j\in\text{ActiveFeatures}}f_{j}$.
@@
  • $p(\neg f_{1},\neg f_{2})=p_{00}$: Probability of neither $f_{1}$ nor $f_{2}$ occurring (e.g., input is 00 or some $f_{k},k\neq 1,2$).

### v6

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across a wide range of tasks, yet our understanding of their internal mechanisms lags behind their capabilities. This gap between performance and interpretability raises concerns about the “black box” nature of these models [rudin2019blackbox]. The field of mechanistic interpretability aims to address this issue by reverse-engineering the internal algorithms of neural networks and performing causal analysis on them [olah2020zoom].
Recent work theorizes that models represent concepts as linear directions in a high-dimensional space, known as the Linear Representation Hypothesis (LRH) [olah2024july, elhage2022toy]. The model is able to represent far more concepts than it has neurons in its hidden space by allowing these concept directions to overlap slighly, known as superposition [elhage2022toy]. Superposition makes it challenging to directly interpret neurons in an LLM, and requires different techniques to extract interpretable feature directions.
As long as the active features in a given LLM activation are sparse and the underlying features follow the LRH, Sparse Autoencoders (SAEs) should be able to recover the true LLM features despite supersition using sparse dictionary learning [olshausen1997sparse]. Indeed, SAEs have shown potential in decomposing the dense, polysemantic activations of LLMs into more “interpretable” latent features [huben2024sparse, bricken2023towards].
However, we show that even if all underlying features are linear and sparsely activating, an SAE will still fail to recover the true underlying features if the features form a hierarchy. Instead, an SAE will learn gerrymandered latents that fail to fire on seemingly arbitrary cases where the latent should fire according to the mainline interpretation of the latent. We refer to this failure as feature absorption.
@@
Feature absorption poses an obstacle to the practical application of SAEs since it suggests SAE latents may be inherently unreliable classifiers. This is particularly important for applications where we need confidence that latents are fully tracking behaviors, such as bias or deceptive behavior. Furthermore, techniques which seek to describe circuits in terms of a sparse combination of latents will also be more difficult in the presence of feature absorption [marks2024feature].
An online explorer for our results can be found at https://feature-absorption.streamlit.app. Code is available at https://github.com/lasr-spelling/sae-spelling.
@@ 2 Background
A linear probe is a simple linear classifier trained on the hidden activations of a neural network, typically using logistic regression (LR) [alain2017understanding].
@@
A k-sparse probe [gurnee2023finding] is a linear probe trained on a sparse subset of $k$ neurons or SAE latents. Training a k-sparse probe first requires selecting the $k$ best neurons or SAE latents that in-aggregate act as a good classifier, and then training a standard linear probe on just those $k$ neurons or latents.
gurnee2023finding proposed several methods of estimating the best $k$ neurons or latents to pick, one of which involves first training a LR probe with a L1 loss term, and selecting the $k$ largest elements by probe weight. When we refer to k-sparse probing in this work, we use this method of selecting $k$ latents.
@@
An SAE consists of an encoder, $W_{enc}$, a decoder, $W_{dec}$, and corresponding biases $b_{enc}$ and $b_{dec}$. The SAE has a nonlinearity, $\sigma$, typically a ReLU (or variant such as JumpReLU [rajamanoharan2024jumping]). Given input activation, $a$, the SAE computes a hidden representation, $f$, and reconstruction, $\hat{a}$:
@@
In an ablation study we examine the downstream causal effect of an SAE latent by computing how patching its activation to 0 changes a downstream metric (e.g. logit difference). A negative ablation effect means intervening on the SAE latent would lower the metric. We follow the work of marks2024feature and provide the algorithm in Appendix A.4.
@@ 3 Toy models of feature absorption
Our initial setup consists of 4 true features, each randomly initialized into orthogonal directions with a 50 dimensional representation vector and unit norm. Each feature fires with magnitude 1.0. Feature $f_{0}$ fires with probability 0.25, and features $f_{1},f_{2}$, and $f_{3}$ fire with probability 0.05. Each SAE training input is created by sampling from these true features and summing the directions of each firing feature. We train a SAE with 4 latents to match the 4 true features using SAELens [bloom2024saetrainingcodebase]. The SAE uses L1 loss with L1 coefficient 3e-5, and learning rate 3e-4. We train on 100M activations.
@@
Figure 2(b) shows the encoder and decoder cosine similarities with the true features in the hierarchical co-occurrence setup. Here, we see a clear example of feature absorption. Latent 0 has learned a perfect representation of feature 0, but the encoder has a hole in its recall. Latent 0 fires if feature 0 is active but not feature 1. This is exactly the sort of gerrymandered feature firing pattern we will see later in real SAEs in Section 5.2 - the encoder has learned to stop the latent firing on specific cases where it looks like it should be firing. In addition, we see that latent 1, which tracks feature 1, has absorbed the feature 0 direction. This results in latent 1 representing a combination of feature 0 and feature 1. We see that the independently firing features 2 and 3 are untouched - the SAE still learns perfect representations of these features. These results are summarized in Table 3. We explore absorption in more toy settings in Appendix A.3.
@@ 4 Experimental setup
In the above prompt, we extract residual stream activations at the _dog token index. These activations are used both for LR probe training and for applying SAEs. We use a train/test split of 80% / 20%, and evaluate only on the test set of the probes, including when running experiments on SAEs. When applying SAEs, we include the SAE error term [marks2024feature] to avoid changing model output.
@@
We use the base Gemma-2-2B model for most of our studies, along with the full set of Gemma Scope residual stream SAEs of width 16k and 65k released by Deepmind [lieberumGemmaScopeOpen2024]. We also evaluate absorption on our own SAEs trained on Qwen2 0.5B [qwen2] and Llama 3.2 1B [dubey2024llama3herdmodels].
@@ 5.1 Do SAEs learn latents that track first letter information?
We also tried using k=1 sparse probing [gurnee2023finding] to identify SAE latents, and found this gives similar results. Further comparison of using k=1 sparse probing vs encoder cosine similarity to identify latents is explored in Appendix A.7.
@@ 5.2 Why do SAE latents underperform?
Figure 6(a) shows a sample prompt containing a series of tokens that start with “S”, and the activations of top SAE latents by ablation score for these tokens. The main “starts with S” latent, 6510, activates on all these tokens except _short. This SAE also has a token-aligned latent, 1085, which activates on variants of the word “short” (“ short”, “SHORT”, etc…). The Neuronpedia dashboard [neuronpedia] for latent 1085 is shown in Appendix A.15. For the token _short, the main “starts with S” latent does not activate but the “short” latent activates instead.
@@
We study the ablation effect of each SAE latent on the _short token, shown in Figure 7(a), and see that latent 1085 has a dramatically larger ablation effect compared with all other SAE latents. This suggests latent 1085 is causally responsible for the model knowing that _short starts with S.
Is it possible that the probe projection is not the causally important component of latent 1085? We conduct another ablation effect experiment, except now we remove the probe direction from latent 1085 via projection before ablation. The results of this experiment are shown in Figure 7(b). After removing the probe component from latent 1085, it no longer has a significant ablation effect. Thus we know the probe projection of latent 1085 is responsible for model behavior.
@@ 5.3 Measuring feature splitting and feature absorption
A key phenomenon identified from previous studies of SAEs is feature-splitting [bricken2023towards], where a feature represented in a single latent in a smaller SAE can split into two or more latents in a larger SAE. During our experiments, we found strong evidence of feature-splitting in the Gemma Scope SAEs.
@@
We measure feature splitting using k-sparse probing [gurnee2023finding] on SAE activations. If increasing the k-sparse probe from $k$ to $k+1$ causes a significant increase in probe F1 score, then the additional SAE latent provides a meaningful signal, and the combination of these $k+1$ latents is likely a feature split. In the example of the uppercase “L” and lowercase “l” split, a k-sparse probe with $k=2$ trained on both these latents should predict “starts with letter L” much better than either latent on its own. Figure 8(a) shows F1 vs K for letters “L” and “N”. The “L” k-sparse probe shows a significant jump in F1 score moving from k=1 to k=2 corresponding to feature splitting, while the F1 score for the “N” k-sparse probe is relatively constant.
@@
If there are more than 200 false negatives per letter, we randomly pick 200 samples to estimate the number of absorptions. We see absorption rate increases with higher sparsity and higher SAE width. Lower L0 likely pushes the SAE to absorb dense features like spelling information, increasing feature sparsity. Feature absorption rate vs L0 for Gemma Scope SAEs layers 0-17 is shown in Figure 9(b). Absorption rate by letter is shown in Appendix A.14. We also train our own set of standard L1 loss SAEs on the first 8 layers of Qwen2 0.5B [qwen2] and Llama 3.2 1B [dubey2024llama3herdmodels], and TopK SAEs [gao2024scaling] on Llama 3.2 1B. In Figure 9(c) we show that absorption occurs in these SAEs as well.
@@ 6 Related work
Probing methods can extract interpretable information from language models, though this does not guarantee the model uses these representations in its computation, and requires labeled data [elazar2021amnesic].
Prior work has shown that many human-interpretable concepts in LLM activations are represented as linear directions in activation space, known as the linear representation hypothesis [elhage2022toy, park2024the]. li2023emergent used non-linear probes to recover board representations from a transformer trained on Othello scripts (“OthelloGPT”). nanda2023actually later showed that linear representations were not only recoverable but also editable.
karvonen2024measuring developed objective metrics for SAE evaluation using Chess and Othello board states, but does not apply these to SAEs trained on LLM activations. Work by olah2024july, kissane2024interpreting, templeton2024scaling noted poor precision/recall of SAE latents compared to known proxies. We extend this by showing how sparsity mediates precision/recall across many Gemma Scope SAEs and offer a possible explanation of low recall due to feature absorption.
engels2025not investigated SAE errors, finding that not all SAE error is linearly decomposable.
@@
Most existing work on SAE interpretability mainly studies max activating examples [huben2024sparse], which may be misleading. There are more rigorous works which only measure precision [bricken2023towards, templeton2024scaling, kissane2024interpreting]. Recent work has briefly explored recall and found it to be worse than expected naively, but this remains poorly understood [olah2024july]. We build on this work by evaluating precision / recall on a large number of SAEs, and offer a partial explanation for lower-than-expected recall of SAE latents in the form of feature absorption.
@@
Feature splitting was first described in bricken2023towards, which noted that different SAE widths and sparsities induce latents of different granularity, with wider SAEs often learning more specific variants of features. bussman2024metasaes find that by training an SAE on the decoder of another SAE, a technique called Meta-SAEs, it is possible to break down a single SAE latent like “Einstein” into subcomponents like “German” and “Physicist” and “starts with E”.
@@ 7 Discussion
The primary goal of future work is to find solutions to feature absorption. We are particularly hopeful that work extending Meta-SAEs [bussman2024metasaes] may solve or mitigate feature absorption. Another possible solution may be attribution dictionary learning [olah2024april]. Finally, structured sparsity techniques such as group lasso [jacob2009group] or hierarchical sparse coding [jenatton2011proximal] may also be a promising direction of future work.
@@ Conclusion
We identify a form of feature splitting we call “feature absorption”, where more specific latents “steal credit” from more general ones. Absorption creates an interpretability illusion, where a seemingly interpretable latent has arbitrary false negatives in its mainline interpretation. Lower recall poses problems for using SAEs for high-stakes classification or finding sparse circuits [marks2024feature], as the number of latents needed to characterize model behavior may be much larger than expected.
@@
This project was produced as part of the LASR Labs research program. DC was supported thanks to EPSRC EP/S021566/1. HB has received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM).
@@ A.1 Glossary of Terms
When inserting a SAE into the computation path of the model, errors in SAE reconstruction will propagate to later parts of the model and can change the model output. We refer to the error as the SAE error term, and corresponds to the difference between the SAE output and the original SAE input activation. marks2024feature introduced the idea of adding this error term back to the SAE output to ensure that the SAE does not change model output.
@@
A platform, https://neuronpedia.org, which hosts feature dashboards for popular SAEs [neuronpedia].
@@
A phenomenon in SAEs introduced by bricken2023towards, where a SAE latent tracks a general feature in a narrow SAE, but splits into multiple more specific SAE latents in a wider SAE. For instance, a latent tracking “starts with L” in a narrow SAE may split into a latent tracking “starts with capital L” and a latent tracking “starts with lowercase L” in a wider SAE.
@@ 1. Preliminaries and Assumptions
Let $\mathcal{D}$ be a dataset. We consider a set of features $\mathcal{F}=\{f_{1},f_{2},\ldots,f_{d}\}$. Each feature $f_{i}\in\mathbb{R}^{k}$ is a vector with unit norm, $\|f_{i}\|_{2}=1$. Features are mutually orthogonal: $f_{i}\cdot f_{j}=\delta_{ij}$ (Kronecker delta), where $\delta_{ij}=1$ if $i=j$ and 0 if $i\neq j$. An activation $h\in\mathbb{R}^{k}$ in the model’s residual stream is a linear combination of active features: $h=\sum_{j\in\text{ActiveFeatures}}f_{j}$.
@@
  • $p(\neg f_{1},\neg f_{2})=p_{00}$: Probability of neither $f_{1}$ nor $f_{2}$ occurring (e.g., input is 0 or some $f_{k},k\neq 1,2$).